# Matched Filtering-Based Channel Estimation for AFDM Systems in Doubly Selective Channels

### v1

@@ III-D Effective SINR Loss
For IDFD channels, $\left|\mathcal{G}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)\right|$ can be expressed as
$\left|\mathcal{G}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)\right|=\left|\frac{\sin\left(\pi\left(\psi-\kappa_{\nu_{i}}\right)\right)}{N\sin\left(\frac{\pi(\psi-\kappa_{\nu_{i}})}{N}\right)}\right|,$ (24)
where $\psi=\left[m-n+2Nc_{1}l_{\tau_{i}}-\bar{k}_{\nu_{i}}\right]_{N}$ is integer. As mentioned before, $\left|h_{w}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)\right|$ has a peak at $\psi=0$, i.e.,
@@
Compared with (23), the SINR loss in the IDFD channels can be expressed as
$L_{\mathrm{SINR}}^{\mathrm{IDFD}}=$ $20\log_{10}\left|\frac{\sin\left(\pi\kappa_{\nu_{i}}\right)}{N\sin\left(\frac{\pi\kappa_{\nu_{i}}}{N}\right)}\right|^{2}\approx 20\log_{10}\left|\frac{\sin\left(\pi\kappa_{\nu_{i}}\right)}{\pi\kappa_{\nu_{i}}}\right|^{2},$ (27)
where ${\pi\kappa_{\nu_{i}}}/{N}$ approaches zero, resulting in $\sin\left({\pi\kappa_{\nu_{i}}}/{N}\right)\approx{\pi\kappa_{\nu_{i}}}/{N}$.
For FDFD channels, similar to the previous derivation, the peak power corresponding to the $i$-th path is
$P_{\mathrm{FDFD}}=\left|h_{i}\right|^{2}\left|\frac{\sin\left(\pi\gamma_{i}\right)}{N\sin\left(\frac{\pi\gamma_{i}}{N}\right)}\right|^{2}.$ (28)
Compared with (23), the SINR loss is
$L_{\mathrm{SINR}}^{\mathrm{FDFD}}=$ $20\log_{10}\left|\frac{\sin\left(\pi\gamma_{i}\right)}{N\sin\left(\frac{\pi\gamma_{i}}{N}\right)}\right|^{2}\approx 20\log_{10}\left|\frac{\sin\left(\pi\gamma_{i}\right)}{\pi\gamma_{i}}\right|^{2}.$ (29)
@@
Fig. 2 shows the effective SINR loss for different fractional delays and fractional Dopplers at $N=64$. The maximum SINR loss is about 7.843 dB. Moreover, according to (27) (29), increasing $N$ cannot effectively reduce the SINR loss. In addition, since $2Nc_{1}\iota_{\tau_{i}}$ can make the fractional part of the delay small enough, we can see that the $\iota_{\tau_{i}}$ has almost no effect on SINR loss when $\kappa_{\nu_{i}}=0$.
Remark 1.
As shown in Fig. 2, the joint effects of fractional delay and fractional Doppler may result in a reduction of $\gamma_{i}$ to alleviate the SINR loss.
IV MF and MF-GFS CE schemes for FDFD channels
@@ IV-A Problem Description
According to (12)-(15) and (64), the system model is
@@
where $\boldsymbol{y}\in\mathbb{C}^{N\times 1}$ and $\boldsymbol{w}\sim\mathcal{C}\mathcal{N}\left(0,\sigma^{2}\boldsymbol{I}_{N}\right)\in\mathbb{C}^{N\times 1}$ are vector forms of $y\left(m\right)$ and $w\left(m\right)$ respectively, $\boldsymbol{I}_{N}$ is the $N$-order identity matrix, vectors $\boldsymbol{l}_{\tau}=\left[l_{\tau_{1}},\cdots,l_{\tau_{P}}\right]^{\mathrm{T}}\in\mathbb{R}^{P\times 1}$, $\boldsymbol{k}_{\nu}=\left[k_{\nu_{1}},\cdots,k_{\nu_{P}}\right]^{\mathrm{T}}\in\mathbb{R}^{P\times 1}$ and $\boldsymbol{h}=\left[h_{1},\cdots,h_{P}\right]^{\mathrm{T}}\in\mathbb{C}^{P\times 1}$ contain the normalized delay, normalized Doppler and channel gain, respectively. ${\boldsymbol{A}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\in\mathbb{C}^{N\times P}$ can be expressed as
@@
Pilots may be inserted into the DAFT domain for efficient CE. Due to the sparsity and compactness of the channel in the DAFT domain, a single pilot can be considered. The arrangement of pilot and data is shown in Fig. 3, which can be expressed as
$x\left(n\right)=\begin{cases}x_{p},&n=n_{p}\\ 0,&n_{p}-Q\leqslant n\leqslant n_{p}+Q,n\neq n_{p}\\ x_{\mathrm{data}},&\mathrm{otherwise}\\ \end{cases},$ (33)
where $x_{p}$ and $x_{\mathrm{data}}$ are single pilot and data, respectively, $n_{p}$ is the pilot index in the DAFT domain grid, $Q\triangleq(l_{\max}+1)(2(k_{\max}+\xi)+1)-1$ due to the structure of $\mathcal{G}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)$ and the analysis in Section III, $Q_{1}=k_{\max}+\xi$, $Q_{2}=Q-Q_{1}$. The DAFT domain grid within the range $n_{p}-Q\leqslant n\leqslant n_{p}+Q,n\neq n_{p}$ remains null as a guard interval to reduce the mutual interference between data and pilot. The two types of guard intervals in Fig. 3 come from pilot and data, respectively.
@@
Based on the pilot pattern and the AFDM symbol structure in Fig. 3, $y\left(m\right)$ in (12) is simplified to
$y\left(m\right)=\sum_{i=1}^{P}{h_{i}\mathcal{G}\left(m,n_{p},k_{\nu_{i}},l_{\tau_{i}}\right)x_{p}}+z\left(n\right),$ (34)
where $n_{p}-Q_{2}\leqslant m\leqslant n_{p}+Q_{1}$, $z\left(n\right)$ is the new noise. The matrix form corresponding to (34) is
@@
IV-B Path Ambiguity in FDFD Channels
Figure 4: Grid arrangement for IDD or IDFD channels.
For both IDD and IDFD channels, the integer parts of the parameters are estimated by
$\hat{\bar{l}}_{\tau_{i}}=\left\lfloor\frac{\hat{\varUpsilon}_{i}}{-2Nc_{1}}\right\rceil,\hat{\bar{k}}_{\nu_{i}}=\hat{\varUpsilon}_{i}+2Nc_{1}\hat{\bar{l}}_{\tau_{i}},$ (37)
where $\hat{\varUpsilon}_{i}$ is the integer part of the estimated equivalent shift $\hat{\chi_{i}}$. That is, the integer delay-Doppler pair $\left(\bar{l}_{\tau},\bar{k}_{\nu}\right)$ for a path uniquely corresponds to a position in the received signal $\boldsymbol{y}_{T}$. This correspondence can be seen in Fig. 4. However, in FDFD channels, (37) is invalid due to the fractional delay, thereby resulting in this one-to-one correspondence no longer holds.
For the FDFD channels, there are additional integer parts. First, $2Nc_{1}\iota_{\tau_{i}}$ leads to a new integer part. In addition, $-\left(2Nc_{1}\iota_{\tau_{i}}-\lfloor 2Nc_{1}\iota_{\tau_{i}}\rceil\right)+\kappa_{\nu_{i}}$ yields another integer part and a final fractional part.
Remark 2.
The extra integer parts may result in the inability to uniquely determine the delay and Doppler of a path. Such a phenomenon is referred to as path ambiguity, which involves delay ambiguity and Doppler ambiguity.
To simplify the explanation, $-2Nc_{1}\iota_{\tau_{i}}+\bar{k}_{\nu_{i}}+\kappa_{\nu_{i}}$ can be regarded as an equivalent Doppler shift of a path, with the range of its integer part represented by
$\mathcal{R}_{\nu}=\begin{cases}\left[-2k_{\max}-\xi-1,k_{\max}\right],&\bar{l}_{\tau_{i}}=0\\ \left[-k_{max},2k_{max}+\xi+1\right],&\bar{l}_{\tau_{i}}=l_{\max}\\ \left[-2k_{\max}-\xi-1,2k_{max}+\xi+1\right],&\mathrm{otherwise}\\ \end{cases}.$ (38)
However, the size of $\mathcal{R}_{\nu}$ is significantly greater than $2Nc_{1}$. Therefore, the grids corresponding to each normalized delay overlap each other. This can be seen in Fig. 5.
Figure 5: Grid arrangement for FDFD channels.
This partial overlap will cause delay ambiguity and Doppler ambiguity. In Fig. 5, there are three arrows representing channel responses, i.e., $E_{0}$, $E_{1}$, and $E_{2}$. Assume that $E_{0}$ corresponds to a path where $l_{\tau_{0}}=\bar{l}_{\tau_{0}}$ and $k_{\nu_{0}}=\bar{k}_{\nu_{0}}+\kappa_{\nu_{0}}<0$. $E_{1}$ corresponds to a path where $l_{\tau_{1}}=\bar{l}_{\tau_{0}}+\iota_{\tau_{1}}$ and $k_{\nu_{1}}=k_{\nu_{0}}$. If $\iota_{\tau_{1}}<0$, we have $-2Nc_{1}\iota_{\tau_{1}}>0$, causing $E_{1}$ to move right relative to $E_{0}$. This leads to Doppler ambiguity. If $-2Nc_{1}\iota_{\tau_{1}}$ is sufficiently large, $E_{1}$ appears in the overlapping region of the grids corresponding to integer delays $1$ and 00, resulting in simultaneous occurrence of both delay ambiguity and Doppler ambiguity. $E_{2}$ corresponds to a path where $l_{\tau_{2}}=\bar{l}_{\tau_{0}}+\iota_{\tau_{2}}$ and $k_{\nu_{2}}=k_{\nu_{0}}$. Similarly, if $\iota_{\tau_{2}}>0$, both delay ambiguity and Doppler ambiguity will also exist simultaneously. Furthermore, due to the additional shift caused by fractional delay, even when an appropriate $\xi$ is set, its ability to mitigate the ICCI between different paths is no longer effective. In addition, $\xi$ may not eliminate path ambiguity. Unfortunately, there is no effective method to solve the path ambiguity problem.
IV-C MF CE for FDFD channels
Figure 6: The proposed pilot pattern in FDFD channel for AFDM systems.
Assume that the wireless channel is a doubly-underspread channel [33], which remains static for a certain period of time. Specifically, the condition for the doubly-underspread channel is given by
$\Delta\tau_{\max}\Delta\nu_{\max}\ll\tau_{\max}\nu_{\max}\ll 1,$ (39)
where $\Delta\tau_{\max}=\frac{\varpi}{c}$, $\Delta\nu_{\max}=2\nu_{\max}\sin\left(\frac{\varrho}{2}\right)$, $c$ is the speed of light, the maximum spatial extension and maximum angular spread of the scatterers are assumed to be $\varpi=30\,\mathrm{m}$ and $\varrho=\frac{\pi}{60}$, respectively. Considering the simulation parameters in Table II of Section VI, we have $\Delta\tau_{\max}\Delta\nu_{\max}=1.0471\times 10^{-5}\ll\tau_{\max}\nu_{\max}=3.12\times 10^{-2}\ll 1$. Therefore, the channel considered is doubly-underspread. Based on $T$ and the stationary time $\frac{1}{\Delta\nu_{\max}}$, the number of AFDM symbols during the stationary time of the channel is $9$.
Therefore, considering the doubly-underspread channel, two AFDM symbols can be used for CE. The key idea is to use the different equivalent shifts of a path in the two AFDM symbols to uniquely determine the delay and Doppler. In addition, the new pilot pattern [^3] An alternative pilot pattern, where $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$ have identical waveform parameters except for $N$, can also achieve the same purpose. in Fig. 6 is proposed to estimate equivalent shifts, where both $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$ have the structure shown in Fig. 3. In addition, for $\boldsymbol{x}_{1}$, the truncated received signal is $\boldsymbol{y}_{\mathrm{T}}$, $c_{1}$ is determined by $\xi$ in (2). For $\boldsymbol{x}_{2}$, similarly, $\boldsymbol{y}_{\mathrm{T}}^{\prime}$, $c_{1}^{\prime}$, and $\xi^{\prime}$ can be defined. Note that setting $\xi\neq\xi^{\prime}$ leads to $c_{1}\neq c_{1}^{\prime}$, thereby resulting in different equivalent shifts.
For the estimation of equivalent shift, let us consider $\boldsymbol{x}_{1}$ as an example. Based on (35), the joint maximum likelihood estimation (JMLE) of $\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu},\boldsymbol{h}\right)$ can be expressed as
@@ IV-C MF CE for FDFD channels
Different columns in $\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ can be approximately orthogonal, making $\boldsymbol{A}_{\mathrm{T}}^{\mathrm{H}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ close enough to the identity matrix. In addition, the larger the $N$ or $\xi$, the better the orthogonality.
@@
Thus, based on the orthogonality, (41) and (42) can be well approximated as
$\left(\hat{\boldsymbol{l}}_{\tau},\hat{\boldsymbol{k}}_{\tau}\right)\approx$ $\mathrm{arg}\max_{\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)}\sum_{i=1}^{P}{\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{y}_{\mathrm{T}}\right|^{2}}.$ (43)
@@
Due to the path ambiguity, $\left(\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ in (45) cannot be estimated accurately. However, the estimation of equivalent shift $\chi_{i}=\varUpsilon_{i}+\gamma_{i}=-2Nc_{1}l_{\tau_{i}}+\small{\frac{\nu_{i}}{\varDelta f}}$ is not affected. Therefore, (45) can be reformulated as
$\hat{\chi}_{i}\approx\mathrm{arg}\max_{\chi_{i}}\left|\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\chi\right)\boldsymbol{y}_{T}\right|^{2},$ (47)
where $\left\{\boldsymbol{\mathcal{T}}\left(\chi\right)\right\}_{m}=\mathcal{J}(m,n_{p},\chi_{i})$, $n_{p}-Q_{2}\leqslant m\leqslant n_{p}+Q_{1}$, $\mathcal{J}(m,n_{p},\chi_{i})$ is defined in (19).
@@
For (47), $\hat{\chi}_{{}_{i}}$ can be obtained by searching for consecutive ${\chi}$. Specifically, since $\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)$ in $\boldsymbol{y}_{\mathrm{T}}$ and $\boldsymbol{\mathcal{T}}\left(\chi\right)$ have the same structure, it can be anticipated that $\left|\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\chi\right)\boldsymbol{y}_{T}\right|$ will achieve its maximum value at $\chi=\chi_{i}$. This proof is similar to Appendix C. However, the complexity of searching for $\hat{\chi}_{i}$ is still high. A simple and effective method is to estimate $\hat{\varUpsilon}_{i}$ using the index $n_{\mathrm{peak}}$ of the maximum peaks in $\boldsymbol{y}_{\mathrm{T}}$, i.e.,
$\hat{\varUpsilon}_{i}=n_{\mathrm{peak}}-n_{p},$ (48)
the fractional $\hat{\gamma}_{i}$ can be obtained by searching within a small range $\left[-0.5,0.5\right]$. For $\hat{\gamma}_{i}$, the continuous range $\left[-0.5,0.5\right]$ can first be uniformly discretized into a grid $\boldsymbol{\varGamma}\in\mathbb{R}^{\left(\rho+1\right)\times 1}$, i.e.,
@@
$\left\{\boldsymbol{\varGamma}\right\}_{\tilde{n}}=-0.5+\frac{1}{\rho}\tilde{n},\tilde{n}\in\left[0,\rho\right],$ (49)
where $\rho$ is the size of the search step. Thus, $\hat{\gamma}_{i}$ is given by
$\hat{\gamma}_{i}=\mathrm{arg}\max_{\gamma\in\boldsymbol{\varGamma}}\left|\boldsymbol{\mathcal{T}}^{\mathrm{H}}\left(\hat{\varUpsilon}_{i}+\gamma\right)\boldsymbol{y}_{\mathrm{T}}\right|^{2}.$ (50)
For $\boldsymbol{x}_{2}$, the equivalent shift $\chi_{i}^{\prime}=\varUpsilon_{i}^{\prime}+\gamma_{i}^{\prime}=-2Nc_{1}^{\prime}l_{\tau_{i}}+\small{\frac{\nu_{i}}{\varDelta f}}$ of the $i$-th path can also be estimated using the above method. Based on the estimated $\hat{\chi}_{i}$ and $\hat{\chi}_{i}^{\prime}$, $\left(\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ can be given by
$\hat{l}_{\tau_{i}}=\frac{\hat{\chi}_{i}^{\prime}-\hat{\chi}_{i}}{2N\left(c_{1}-c_{1}^{\prime}\right)},$ (51)
$\hat{k}_{\nu_{i}}=\frac{\hat{\nu}_{i}}{\varDelta f}=\hat{\chi}_{i}+2Nc_{1}\hat{l}_{\tau_{i}}.$ (52)
@@
Noted that $\left(\hat{h}_{i},\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ in (46), (48)-(52) represent the estimation results for single path. For multiple paths, an iterative estimation scheme is required. Specifically, the signal corresponding to $\left(\hat{l}_{\tau}^{\left(t\right)},\hat{k}_{\nu}^{\left(t\right)},\hat{h}^{\left(t\right)}\right)$ in the $\left(t\right)$-th iteration is
@@
For $\boldsymbol{y}_{\mathrm{T}}^{\prime}$, the same process is performed. The iteration terminates if the maximum number $T_{\mathrm{iter}}$ is reached or $\left|\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t+1\right)}\right\|-\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t\right)}\right\|\right|/\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t\right)}\right\|\leqslant\sigma$, where the preset threshold $\sigma>0$. The MF CE scheme is summarized in Algorithm 1.
Algorithm 1 Proposed MF CE scheme for FDFD channels
0:  Received signal $\boldsymbol{y}_{\mathrm{T}}$ and $\boldsymbol{y}_{\mathrm{T}}^{\prime}$. Initialization: ${\boldsymbol{y}_{\boldsymbol{T}}^{\prime}}^{\left(t\right)}=\boldsymbol{y}_{\mathrm{T}}^{\prime}$, $\boldsymbol{y}_{\mathrm{T}}^{\left(t\right)}=\boldsymbol{y}_{\mathrm{T}}$, the maximum number of iterations $T_{\mathrm{iter}}$, threshold $\sigma$, $l_{\max}$, $k_{\max}$, $\hat{\boldsymbol{l}}_{\tau}=\left[\right],\hat{\boldsymbol{k}}_{\nu}=\left[\right],\hat{\boldsymbol{h}}=\left[\right]$, $t=1$.

### v3

@@ III-E Effective SINR Loss
For IDFD channels, based on subsection III-B-2), we have
@@
Compared with (29), the SINR loss in the IDFD channels is
$L_{\mathrm{SINR}}^{\mathrm{IDFD}}=$ $2,\displaystyle 0\log_{10}\left|\frac{\sin\left(\pi\kappa_{\nu_{i}}\right)}{N\sin\left(\frac{\pi\kappa_{\nu_{i}}}{N}\right)}\right|^{2}\approx 0\log_{10}\left|\frac{\sin\left(\pi\kappa_{\nu_{i}}\right)}{\pi\kappa_{\nu_{i}}}\right|^{2}$, (32)
where ${\pi\kappa_{\nu_{i}}}/{N}$ approaches zero, resulting in $\sin\left({\pi\kappa_{\nu_{i}}}/{N}\right)\approx{\pi\kappa_{\nu_{i}}}/{N}$. In general, $N\gg\pi k_{\nu_{i}}$ is satisfied, therefore this approximation holds. Moreover, according to (32), increasing $N$ cannot effectively reduce the SINR loss.
For FDFD channels, based on (22)-(23), the peak value $\mathcal{F}_{\max}$ of $\left|\mathcal{F}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)\right|$ can be approximated as
$\mathcal{F}_{\max}\approx\left|\frac{1}{N}\sum_{\bar{n}=0}^{N-1}{e^{\small{\frac{-j2\pi}{N}}\eta_{m_{\mathrm{m}-\mathrm{peak}}^{\left(i\right)},n}^{\left(i\right)}\bar{n}}}\right|\approx\left|\frac{\sin\left(\pi\kappa_{\nu_{i}}\right)}{\pi\kappa_{\nu_{i}}}\right|.$ (33)
Therefore, increasing $N$ does not reduce the SINR loss in the FDFD channel. The SINR loss in the FDFD channels is
[…]
@@
Based on (34), Fig. 2 shows the effective SINR loss for different fractional delays and fractional Dopplers at $N=64$. Increasing $\iota_{\tau_{i}}$ or $\kappa_{\nu_{i}}$ results in more severe SINR loss. The maximum SINR loss is about 11.788 dB. Moreover, according to (32) and (33), increasing $N$ cannot effectively reduce the SINR loss.
IV MF and MF-GFS CE Schemes
@@ IV-A Problem Description
According to (13)-(17), the system model is
@@
where $\boldsymbol{y}\in\mathbb{C}^{N\times 1}$ and $\boldsymbol{w}\sim\mathcal{C}\mathcal{N}\left(0,\sigma^{2}\boldsymbol{I}_{N}\right)\in\mathbb{C}^{N\times 1}$ are vector forms of $y\left(m\right)$ and $w\left(m\right)$ respectively, vectors $\boldsymbol{l}_{\tau}=\left[l_{\tau_{1}},\cdots,l_{\tau_{P}}\right]^{\mathrm{T}}\in\mathbb{R}^{P\times 1}$, $\boldsymbol{k}_{\nu}=\left[k_{\nu_{1}},\cdots,k_{\nu_{P}}\right]^{\mathrm{T}}\in\mathbb{R}^{P\times 1}$ and $\boldsymbol{h}=\left[h_{1},\cdots,h_{P}\right]^{\mathrm{T}}\in\mathbb{C}^{P\times 1}$ contain the normalized delay, normalized Doppler and channel gain, respectively. ${\boldsymbol{A}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\in\mathbb{C}^{N\times P}$ can be expressed as
@@
1) Pilot Pattern: Pilots may be inserted into the DAFT domain for efficient CE. Due to the sparsity and compactness of the channel in the DAFT domain, a single pilot can be considered. The arrangement of pilot and data is shown in Fig. 3, which can be expressed as
[…]
where $x_{p}$ and $x_{\mathrm{data}}$ are single pilot and data, respectively, $n_{p}$ is the pilot index in the DAFT domain grid, $Q\triangleq(l_{\max}+1)(2(k_{\max}+\xi)+1)-1$ due to the structure of $\mathcal{G}\left(m,n,k_{\nu_{i}},l_{\tau_{i}}\right)$ and the analysis in Section III. The DAFT domain grid within the range $n_{p}-Q\leqslant n\leqslant n_{p}+Q,n\neq n_{p}$ remains null as a guard interval to reduce the mutual interference between data and pilot. Since $\bar{l}_{\tau_{i}}$ in the equivalent shift $\chi_{i}=-2Nc_{1}\bar{l}_{\tau_{i}}+\bar{k}_{\nu_{i}}$ is non-negative, the guard intervals on the two sides of the pilot have different lengths. Thus, the right and left guard intervals are of lengths $Q_{1}=k_{\max}+\xi$ and $Q_{2}=Q-Q_{1}$, respectively. In addition, the two colored guard intervals in Fig. 3 come from pilot and data, respectively.
@@
2) CE Model: Since the main pilot response in $\boldsymbol{y}$ is distributed in the region for CE, truncating the $\boldsymbol{y}$ is intended to reduce complexity. In particular, we employ the received signals $y\left[m\right]$ in the range of $n_{p}-Q_{2}\leqslant m\leqslant n_{p}+Q_{1}$ to estimate the channel. Therefore, $y\left[m\right]$ in (13) is simplified to
$y\left[m\right]=\sum_{i=1}^{P}{h_{i}\mathcal{G}\left(m,n_{p},k_{\nu_{i}},l_{\tau_{i}}\right)x_{p}}+z\left[n\right],$ (39)
where $z\left[n\right]$ is the new noise. The matrix form corresponding to (39) is
@@
3) Decoupling Multipath Estimation: Based on (40), the joint maximum likelihood estimation (JMLE) of $\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu},\boldsymbol{h}\right)$ can be expressed as
@@
For any $i\neq j$, $\left\|\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\right\|^{2}=1$,
$\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|\leqslant\epsilon,$ (45)
where $\epsilon=\underset{i\neq j}{\max}\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{j}},k_{\nu_{j}}\right)\right|$.
Remark 1.
Since $\epsilon\ll 1$, $\boldsymbol{A}_{\mathrm{T}}^{\mathrm{H}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)\boldsymbol{A}_{\mathrm{T}}\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)$ is approximately the identity matrix $\boldsymbol{I}_{P}$.
@@
Thus, (43) and (44) can be well approximated as
$\left(\hat{\boldsymbol{l}}_{\tau},\hat{\boldsymbol{k}}_{\tau}\right)\approx$ $\mathrm{arg}\max_{\left(\boldsymbol{l}_{\tau},\boldsymbol{k}_{\nu}\right)}\sum_{i=1}^{P}{\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)\boldsymbol{y}_{\mathrm{T}}\right|^{2}},$ (46)
@@
IV-B MF CE for FDFD channels
@@
For (48), $\left(\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ can be obtained by searching for consecutive $\left(l_{\tau_{i}},k_{\nu_{i}}\right)$. Specifically, since $\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau_{i}},k_{\nu_{i}}\right)$ in $\boldsymbol{y}_{\mathrm{T}}$ and $\boldsymbol{a}_{\mathrm{T}}\left(l_{\tau},k_{\nu}\right)$ have the same structure, it can be anticipated that $\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(l_{\tau},k_{\nu}\right)\boldsymbol{y}_{T}\right|$ achieves its maximum value at $\left(l_{\tau},k_{\nu}\right)=\left(l_{\tau_{i}},k_{\nu_{i}}\right)$. However, the complexity of searching for $\left(\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ in space $\left\{\left[0,l_{\max}\right]\times\left[-k_{\max},k_{\max}\right]\right\}$ is still high. A simple and effective method is to estimate $\left(\bar{l}_{\tau_{i}},\bar{k}_{\nu_{i}}\right)$ of the path based on the maximum peaks in the received signal $\boldsymbol{y}_{\mathrm{T}}$, i.e.,
$\left.\hat{\bar{l}}_{\tau_{i}}=\left.\lfloor\frac{\hat{\chi}_{i}}{-2Nc_{1}}\right.\right.\rceil,\hat{\bar{k}}_{\nu_{i}}=\hat{\chi}_{i}+2Nc_{1}\hat{\bar{l}}_{\tau_{i}},$ (50)
where $\hat{\chi}_{i}=m_{\max}-n_{p}$, $m_{\max}=\mathrm{arg}\underset{m}{\max}\left|\left\{\boldsymbol{y}_{\mathrm{T}}\right\}_{m}\right|$.
The fractional $\left(\hat{\iota}_{\tau_{i}},\hat{\kappa}_{\nu_{i}}\right)$ can be obtained by a global search within a small continuous region $\left\{\left[-0.5,0.5\right]\times\left[-0.5,0.5\right]\right\}$. Therefore, $\left(\hat{\iota}_{\tau_{i}},\hat{\kappa}_{\nu_{i}}\right)$ can be expressed as
$\left(\hat{\iota}_{\tau_{i}},\hat{\kappa}_{\nu_{i}}\right)=\mathrm{arg}\max_{\left(\iota_{\tau},\kappa_{\nu}\right)\in\boldsymbol{\varOmega}}\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}}+\iota_{\tau},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{y}_{T}\right|^{2},$ (51)
@@
where $\boldsymbol{\varOmega}=\boldsymbol{\varGamma}\times\boldsymbol{\varGamma}\in\mathbb{R}^{\left(\rho+1\right)\times\left(\rho+1\right)}$, $\varGamma_{\tilde{n}}=-0.5+\frac{1}{\rho}\tilde{n},\tilde{n}\in\left[0,\rho\right]$, $\rho$ is the size of the search step.
To further reduce the complexity, a decoupled DD estimation can be considered, i.e.,
$\hat{\iota}_{\tau_{i}}=\mathrm{arg}\max_{\iota_{\tau}\in\boldsymbol{\varGamma}}\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}}+\iota_{\tau},\hat{\bar{k}}_{\nu_{i}}\right)\boldsymbol{y}_{T}\right|^{2},$ (52)
$\hat{\kappa}_{\nu_{i}}=\mathrm{arg}\max_{\kappa_{\nu}\in\boldsymbol{\varGamma}}\left|\boldsymbol{a}_{\mathrm{T}}^{\mathrm{H}}\left(\hat{\bar{l}}_{\tau_{i}}+\hat{\iota}_{\tau_{i}},\hat{\bar{k}}_{\nu_{i}}+\kappa_{\nu}\right)\boldsymbol{y}_{T}\right|^{2}.$ (53)
Compared with the joint DD estimation in (51), which requires $\left(\rho+1\right)^{2}$ searches, the decoupled DD estimation in (52)–(53) involves only $2\left(\rho+1\right)$ searches. To distinguish between the joint DD estimation and the decoupled DD estimation, they are referred to as MF-JE and MF-DE, respectively.
@@
Noted that $\left(\hat{h}_{i},\hat{l}_{\tau_{i}},\hat{k}_{\nu_{i}}\right)$ in (49)-(53) represent the estimation results for single path. For multiple paths, an iterative estimation scheme is required. Specifically, the signal corresponding to $\left(\hat{l}_{\tau}^{\left(t\right)},\hat{k}_{\nu}^{\left(t\right)},\hat{h}^{\left(t\right)}\right)$ in the $\left(t\right)$-th iteration is
@@
The iteration terminates if the maximum number $T_{\mathrm{iter}}$ is reached or $\left|\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t+1\right)}\right\|-\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t\right)}\right\|\right|/\left\|\boldsymbol{y}_{\mathrm{T}}^{\left(t\right)}\right\|\leqslant\sigma$, where the preset threshold $\sigma>0$. The MF with joint DD estimation (MF-JE) CE scheme is summarized in Algorithm 1.
Algorithm 1 Proposed MF-JE and MF-DE CE scheme for FDFD channels
[…]